OpenStudy (anonymous):

What is the equation of the line that contains the point (–5, –1) and has a slope of 4? Write in slope-intercept form. A. y = 4x – 19 B. y = 4x + 19 C. y = 4x – 1 D. y = 4x + 1

OpenStudy (anonymous):

you can try every answer.... for example a. -1=4*-5-19 what idnt equal

OpenStudy (anonymous):

y = m x+b The numberic values for x, y and m is given in the problem statement. Plug them in the general line equation above and solve for b.

OpenStudy (anonymous):

so the answer is A

OpenStudy (anonymous):

b is the answe

OpenStudy (anonymous):

can u help me with another one

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is the equation, in slope-intercept form, of the line that has a slope of –3 and passes through the point (0, 5)?

OpenStudy (anonymous):

y=ax+b, a is the slope, b is intersection with x ray, so y=-3x+5

OpenStudy (anonymous):

another one?

OpenStudy (anonymous):

go to school :P

OpenStudy (anonymous):

lol please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the y-intercept of the line that has a slope of -1/2and passes through the point (2, 3)?

OpenStudy (anonymous):

y=-1/2x+?

OpenStudy (anonymous):

i told you what in y=ax+b is slope, its a..., if (2,3) is point of y=ax+b then 3=2a+b. so because a=-1/2 3=-1+b ->b=4 so ->y=-1/2*a+4, y intercept is 4

OpenStudy (anonymous):

another one?

OpenStudy (anonymous):

Using point-slope form, write the equation of the line that passes through the point (-4, 12) and has a slope of -3/4

OpenStudy (anonymous):

Refer to the two line solution from Mathematica.

OpenStudy (anonymous):

@robtobey thanks but i already got the answers but can help me out with this one

OpenStudy (anonymous):

Using point-slope form, write the equation of the line that passes through the point (-4, 12) and has a slope of -3/4

OpenStudy (anonymous):

See the second attachment below. It's the same process.

OpenStudy (anonymous):

(y+12)=-3/4(x-4)

OpenStudy (anonymous):

@robtobey